OpenStudy (anonymous):

Check my answer please! Will grant a medal Decide whether the equation has one solution, no solution, or infinitely many solutions. 3(y - 3) = 2y - 9 + y My answer: **Infinitely many solutions**

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

3(y-3)=2y 3y-9=2y y=9 2y=9+y y=9

OpenStudy (anonymous):

That is not the whole equation though. This is: 3(y - 3) = 2y - 9 + y

OpenStudy (anonymous):

aha

OpenStudy (anonymous):

you'r correct

OpenStudy (anonymous):

Thanks